OpenStudy (anonymous):

Simplify. –12 – 4.8 ∙ (–2 + –8) A. –168 B. 36 C. 72 D. 168

OpenStudy (brittanie1002):

D

OpenStudy (anonymous):

can u help with 2 more

OpenStudy (anonymous):

Evaluate the expression when x = –15. –4x – 11 A. –71 B. –49 C. 49 D. 71

OpenStudy (brittanie1002):

I CAN TRY

OpenStudy (texaschic101):

-12 - 4.8 * (-2 + (-8) = -12 - 4.8 * (-10) = -12 + 48 = 36 -4x - 11....when x = -15 -4(-15) - 11= 60 - 11 = 49

OpenStudy (anonymous):

Evaluate the expression when p = 9.3 and q = –2.1. p – 3q A. 15.6 B. 3.0 C. –13.2 D. –15.6

OpenStudy (anonymous):

9.3-3-2.1

OpenStudy (texaschic101):

p - 3q 9.3 - 3(-2.1).....do the multiplication first and remember that a neg times a neg is a positive. now what do you get ?

OpenStudy (texaschic101):

what is -3 * -2.1 ??

OpenStudy (anonymous):

0.9?

OpenStudy (anonymous):

or -5.1

OpenStudy (texaschic101):

we are multiplying.....what is -3 x -2.1....and remember that a neg times a neg is a positive

OpenStudy (anonymous):

6.2

OpenStudy (texaschic101):

close...it is 6.3 so now we have : 9.3 + 6.3....so just add

OpenStudy (anonymous):

15.6

OpenStudy (texaschic101):

correct :)

OpenStudy (anonymous):

yay i got an 80 thanks

OpenStudy (texaschic101):

anytime :)

OpenStudy (anonymous):

still passed

OpenStudy (anonymous):

:)